# Near instantaneous O(1) Analog Solver Circuit for Linear Symmetric Positive-Definite Systems

### v1

@@ IV-B System Testing
where substituting $D_{ii}$ from Equation 22 yields:
@@
This condition indicates that the system $(A-K_{s})$ must be diagonally dominant. Consequently, solving a diagonally dominant symmetric matrix with this design achieves a complexity $O(1)$, which is the theoretical maximum speed attainable for the solution process.
IV-B1 Complexity Studies
@@
Figure 12: Analysis of 1200 random systems (density=1) using the proposed design. Plot (a) shows the maximum conductance in the transformed system of each system. Plot (b) shows the maximum convergence time for each system
To evaluate the performance of the proposed design in terms of error and speed within the domain of symmetric positive definite systems that are not diagonally dominant, both operating point and transient analyses were conducted on randomly generated systems in this domain. For each randomly generated conductance matrix $A$, unknowns $x_{i}$ were randomly assigned values within the range [-0.5V,0.5V], and the corresponding vector $b$ was calculated by $b=Ax$.
[…]
The proposed design demonstrated a remarkable improvement in convergence time, achieving at least a 100x speedup compared to the preliminary design. Systems that required milliseconds to solve using the preliminary design were solved in only tens of microseconds with the proposed design. Furthermore, the variation in convergence time across simulations was significantly reduced, highlighting the efficiency and consistency of the proposed design. Although the results in Figure 12 suggest a linear relationship between the settling/convergence time of the systems and the number of unknowns, the generated systems also exhibit a linear dependency between the number of unknowns and the maximum conductance in the transformed system. This indicates that the settling time may be influenced by either the number of unknowns or the maximum conductance in the transformed system.
@@
Figure 13: Analysis of 900 systems with maximum conductance within 10% of 800 $\mu S$ (density=1) using the proposed design. Plot (a) shows the maximum error between the converged solution and the theoretical values of each system. Plot (b) shows the maximum convergence time for each system
@@
Figure 14: Analysis of 1200 systems with maximum conductance within 10% of 550 $\mu S$ (density=0.5) using the proposed design. Plot (a) shows the maximum error between the converged solution and the theoretical values of each system. Plot (b) shows the maximum convergence time for each system
In the proposed design, the resistors connecting the nodes to the ground are minimal or nonexistent, meaning that the maximum conductance in the transformed system corresponds to the maximum conductance of the resistors connecting the nodes representing the unknowns. This is reflected in the off-diagonal terms of the transformed matrix, which are equal to the off-diagonal terms of the original $A$ matrix. Derived from Equations 22 and 24, the diagonal elements of the sub-matrix $K_{B}$, part of the off-diagonal terms of the transformed system, are computed assuming the diagonals of $A$ are positive as:
$K_{Bii}=-\frac{1}{2}\left(A_{ii}-K_{sii}-\sum_{j=1,j\neq i}^{n}|A_{ji}|\right)$ (26)
These diagonal terms represent the deviation of the matrix $(A-K_{s})$ from diagonal dominance. The complexity studies demonstrate that the speed and error in solving symmetric positive definite systems that are not diagonally dominant using the proposed design depend on key matrix properties: the eigenvalues, the maximum absolute values of the off-diagonal terms, and the deviation from diagonal dominance. Importantly, these factors show that the performance does not depend on the size of the system being solved.
IV-B2 Operational Amplifier Specifications
@@
The accuracy and speed of the proposed design are influenced by the specifications of the operational amplifiers used. To examine the impact of amplifier choice, the same set of random matrices used in the preliminary design and test 1 of the complexity studies were simulated using three different amplifiers. The first amplifier, AD712, was used in all previous simulations. The second amplifier, LTC2050, features a low input offset voltage to minimize errors in the converged solutions. The third amplifier, LTC6268, has relatively high slew rate and gain bandwidth product to reduce settling time. The main specifications of these amplifiers are summarized in Table I.
@@
The simulation results, presented in Figure 15, reveal a tradeoff between speed and accuracy. Using the LTC2050 amplifier reduces the error in the converged solution by an order of magnitude compared to the AD712 but increases the settling time by approximately 1.6 times. Conversely, the LTC6268 amplifier significantly speeds up the system, achieving a two-orders-of-magnitude reduction in settling time, but with an increase in error by an order of magnitude. This tradeoff highlights that a lower input offset voltage results in higher accuracy in the converged solution, while a higher gain bandwidth product leads to faster system speed.
@@
IV-B3 System Scaling
The motivation behind the proposed design is to optimize performance in terms of speed, accuracy, and power consumption. A detailed analysis of the system power consumption is presented in Section IV-B4. One approach to reducing power consumption is to scale the system of equations such that the conductance values of both the left-hand and right-hand side models are proportionally scaled down. When both models are scaled by the same factor, denoted as $\alpha$, the resulting solution for the unknown vector $x$ remains unchanged. The scaled system is expressed as:
$\alpha\begin{bmatrix}K_{A}&K_{B}\\ K_{B}&K_{A}\end{bmatrix}\begin{Bmatrix}x\\ -x\end{Bmatrix}=\alpha\begin{Bmatrix}b-K_{s}x\\ -b-K_{s}(-x)\end{Bmatrix}$ (27)
To study the effect of the scaling factor $\alpha$ on the error and settling time, random symmetric positive definite matrices that are not diagonally dominant were generated and simulated for different values of $\alpha$. Figure 16 presents results for two systems: a 5x5 system and a 20x20 system. The simulations consistently show that reducing $\alpha$ decreases both the error in the converged solution and the convergence time. Therefore, scaling the system down is highly advantageous, as it enhances accuracy, speeds up the solution process, reduces power consumption, and adjusts the conductance values of resistors to a commercially common range. Additionally, operating the design in the low conductance (high resistance) range minimizes errors from parasitic resistance by reducing the relative impact of parasitic resistances compared to the circuit resistors.
Figure 16: Analysis of two random systems using the proposed design for different values of system scaling $\alpha$. Plot (a) shows the maximum error between the converged solution and the theoretical values of each system. Plot (b) shows the maximum convergence time for each system.
IV-B4 Power Consumption
The power consumption of the proposed design is analyzed in this section, assuming the circuit operates at the converged solution of the system. The analogy between resistors and 1D springs is employed to calculate the system power. In this analogy, the power equation for a resistor is equal to the energy equation for a spring, where the conductance of the resistor represents the stiffness of the spring, and the voltage difference across the resistor corresponds to the displacement difference between the spring’s end nodes. The total energy in a spring system is calculated using the conservation of energy between the external energy (from the applied forces or the right-hand side) and the internal energy stored in the springs.
Applying this conservation principle to the proposed design, the total power $P_{k^{+}}$ of the left-hand-side model, assuming only positive resistors, is expressed as:
$P_{k^{+}}=\sum_{k=1}^{n_{el}}k_{k}(x_{i}-x_{j})^{2}=x_{2n}^{T}K_{2n}x_{2n},$ (28)
where $n_{el}$ is the number of resistors in the left-hand-side, $k_{k}$ is the conductance of the individual resistors, $x_{i}$ and $x_{j}$ are the voltages at the two ends of the resistor, and $K_{2n}$ is the transformed conductance matrix of the proposed design.
Since this equation assumes all conductance values are positive, a correction term is necessary to account for negative resistance. According to the choice of the $D$ matrix presented in Section IV-A3, negative resistance exist only between nodes $i$ and $n+i$, where the voltage difference is $2x_{i}$. The power correction term $P_{k^{-}}$, accounting for two resistors ($R_{pot1}$$R_{pot2}$) with the same conductance in the negative resistance circuit, is:
$P_{k^{-}}=6x^{T}\left(K_{B}+|K_{B}|\right)x.$ (29)
The power of the remaining components of the left-hand-side model of the system is derived from analyzing the circuit in Figure 3 with buffers, where nodes $i$ and $j$ in this case are nodes $i$ and $n+i$. The output voltage at the buffers are $x_{i}$ and $x_{i}$, while the output voltage at the opamps output are $3x_{i}$ and $3x_{i}$. Thus, the power consumption $P_{R}$ of the resistors with conductance $k_{R}$ used to control the gain of the circuit can be expressed as follows:
$P_{R}=4k_{R}x^{T}x$ (30)
Combining these equations and simplifying Equation 28 using Equations 14 and 18, and adding the power consumption from the right-hand side $(2x^{T}K_{s}x)$, the total system power $P_{sys}$ is:
$P_{sys}=P_{amp}+P_{sw}\\ +4k_{R}x^{T}x+6x^{T}(K_{B}+|K_{B}|)x+2x^{T}Ax,$ (31)
where $P_{amp}$ and $P_{sw}$ are the total power consumed by amplifiers and switches, respectively. This equation highlights that power consumption decreases for systems with lower conductance, making it advantageous to scale down the system. Additionally, the power consumption of the operational amplifiers can be further reduced by lowering the supply voltage to the amplifiers, ensuring that it remains sufficient to accurately output three times the maximum expected voltage in the solution vector.
@@ V Discussion
  • Diagonally Dominant Systems: For these systems, the solution is achieved with $O(1)$ complexity, irrespective of the matrix size or specific properties, since the circuit consists exclusively of resistors. This enables the system to reach the theoretical maximum speed.
  • Non-Diagonally Dominant Systems: The accuracy and speed of the solution depend on the specifications of the opamps used and the system matrix properties such as eigenvalues, the maximum off-diagonal term, and deviation from diagonal dominance. However, for systems with similar eigenvalues and matrix properties, the solution still achieves $O(1)$ complexity.
While the proposed design excels in speed and accuracy, it faces a computational setback in calculating the sum of the absolute values of the columns, which has a complexity of $O(n^{2})$. To mitigate this issue, several solutions are proposed.
  • System Assembly: In applications such as the Finite Element Method, calculating the sum is computationally inexpensive when calculated as part of the construction and assembly of the linear system of equations
  • Parallel Resistors: Doubling the number of variable resistors and constructing the variable resistors representing the diagonals of the submatrix $K_{B}$ using individual variable resistors in parallel, each with conductance corresponding to individual elements in the corresponding column.
  • Matrix Vector Multiplication (MVM): Utilizing an analog MVM circuit to multiply the matrix by a vector of ones, reducing the complexity of this operation to $O(1)$
@@ VI Conclusion
The presented research work introduces a fundamentally new class of analog compute-in-memory (CIM) solvers, resistive network mapping, that directly maps the given SPD system of equations into a resistive network. This mapping is only realized after leveraging the noninverting operational amplifier configuration with some modifications to map negative conductance. Another key aspect of the architecture is the proposed transformation of the original $n\times n$ system (up to $(n^{2}-n)/2$ negative conductance values) into an equivalent $2n\times 2n$ system (up to $n$ negative conductance values). Mapping and solving this transformed system reduces the needed component count by more than 70%, consumes less power, and speeds up the computation by orders of magnitude without additional computational overhead.
@@
For the subclass of symmetric diagonally dominant (SDD) matrices, mapping the transformed system results in a purely passive resistive network whose nodal voltages settle in effectively zero time. This yields true $O(1)$ complexity where the solution speed reaches the theoretical speed limit for analog-CIM architectures and outperforms the state-of-the-art analog solvers, whose settling times are limited by RC constants and feedback loop dynamics, by orders of magnitude. Compared to conventional digital algorithms (even highly optimized sparse solvers), the passive resistive network achieves multiple orders of magnitude speedups for large SDD systems, while consuming a fraction of the energy.
More generally, for any SPD system, the solution speed remains independent of the number of unknowns and depends only on other matrix properties such as eigenvalues, maximum off-diagonal term, and deviation from diagonal dominance. The accuracy of the proposed design is limited by the specifications of the variable resistors and opamps used, which will keep improving with the rapid improvements in device technology. Numerical convergence studies of non-diagonally dominant SPD systems show four significant figure accuracy under idealized conditions. Nonetheless, the sensitivity of solution to component tolerances and environmental perturbations warrants further investigations.
@@
A notable advantage of the proposed design is its compatibility with the cross-point architecture, commonly used in resistive memory arrays. This compatibility facilitates in-memory computations, enabling multiple operations such as MVM within the same framework.
@@
In summary, Resistive Network Mapping presents a compelling paradigm shift in computational linear algebra by eliminating the mathematical process of solving linear systems of equations. The proposed design offers substantial improvements in speed and efficiency, particularly for SPD systems, such as those arising from Finite Element Analysis in Computational Solid Mechanics. Future work could explore extending this design to a broader range of linear systems and developing more robust solutions for handling non-positive definite matrices efficiently.

### v2

@@ IV-C3 Domain of Positive Resistors
where substituting $D_{ii}$ from Equation (31) yields:
@@
This condition indicates that the system $(A-K_{s})$ must be diagonally dominant. The settling time of resistive networks is governed by the RC constants arising from parasitic capacitance in the network. However, the RC delay in resistive memristor arrays is in the order of tens of picoseconds [11] and therefore can be neglected. Consequently, solving a symmetric diagonally dominant (SDD) matrix with this design achieves a complexity $O(1)$, which is the theoretical maximum speed attainable for the solution process.
@@
Figure 13: Analysis of 1200 random dense systems (eigenvalues: [10 1000] $\mu S$) using the proposed design. Plot (a) shows the maximum conductance in the transformed system of each system. Plot (b) shows the maximum convergence time for each system
IV-C4 Bound Analysis
To gain deeper insight into the dynamics of the proposed circuit, upper and lower bounds on its settling time are derived. Since the governing differential equation is an overdamped second-order equation, adding a diagonal damping term increases the settling time and thus provides a valid upper bound. By adding the term $\left(2\tau K_{N}\dfrac{du}{dt}\right)$ to the governing Equation (44), the differential equation reduces to the following
$2\tau^{2}\frac{d^{2}u}{dt^{2}}+3\tau\frac{du}{dt}=-M^{-1}(Ku-f)$ (55)
This upper-bound system remains a second-order differential equation that solves the preconditioned linear system $M^{-1}Ku=M^{-1}f$.
A lower bound on the settling time can be obtained by neglecting the delay introduced by the buffer OAs inside the negative resistance circuit. Applying Equation (18) for the transformed system and using the voltage relationship $u_{j}=-u_{n+j}$, the differential equation in vector form can be written as
$2\tau\frac{du^{\prime}}{dt}=3u-u^{\prime}$ (56)
Substituting Equations (41) and (43) in Equation (56) results in the following lower bound:
$2\tau\frac{du}{dt}=-M^{-1}(Ku-f)$ (57)
This lower bound is a first-order differential equation that also solves the same preconditioned system of equations $M^{-1}Ku=M^{-1}f$.
Both bounds demonstrate that the circuit dynamics are governed not by the original system $Ku=f$, but rather by its preconditioned form with $M^{-1}$ as the preconditioner. For matrices close to diagonal dominance, $M^{-1}K$ is nearly the identity matrix, resulting in a more well-conditioned system that reduces error amplification.
IV-D Complexity Studies for SPD Systems
To evaluate the performance of the proposed design, complexity studies are performed for different classes of systems and compared to the feedback circuit in [41].
IV-D1 Spectrum-Constrained Random Matrices
@@
Figure 14: Analysis of 900 systems (eigenvalues: [10 1000] $\mu S$) with maximum DD-deviation $R_{max}$ within 10% of 800 $\mu S$ (density=1) using the proposed design. Plot (a) shows the maximum error between the converged solution and the theoretical values of each system. Plot (b) shows the maximum convergence time for each system
The complexity studies in this section follow the same methodology used for the preliminary design studies. The objective is to validate that the computational complexity of the proposed design does not depend on the size of the system, but on matrix properties,specifically, the smallest eigenvalue and DD deviation.
Random SPD but non-SDD matrices of various sizes were generated using MATLAB and then simulated in LTspice v.17.1.8 for both operating point and transient analyses. The OA model used is the manufacturer-supplied SPICE model for the AD712 which incorporates several nonideal behaviors, including finite bandwidth, finite slew rate, input offset voltage, bias currents, nonlinear output swing limits, and internal resistances. This more realistic OA model allows the study to capture circuit behavior beyond the simplified single-pole approximation.
The random SPD systems generated in these studies eigenvalues ranges between and includes 10 $\mu S$ and 1000 $\mu S$. For each randomly generated conductance matrix $A$, unknowns $x_{i}$ were randomly assigned values within the range [-0.5V,0.5V], and the corresponding vector $b$ was calculated by $b=Ax$. In the transient simulations, the supply voltage was modeled as a step function that changes from zero to $x_{s}^{+}$ or $x_{s}^{-}$ at 1 $\mu s$. This setup was used to analyze the settling time of the system and the accuracy of the solution for the unknown vector $x$.
Figure 15: Analysis of 1200 systems (eigenvalues: [10 1000] $\mu S$) with maximum DD-deviation within 10% of 550 $\mu S$ (density=0.5) using the proposed design. Plot (a) shows the maximum error between the converged solution and the theoretical values of each system. Plot (b) shows the maximum convergence time for each system
@@
Figure 16: Analysis of 1200 systems (eigenvalues: [10 1000] $\mu S$) with maximum DD-deviation within 10% of 550 $\mu S$ (density=0.5) using the proposed design and the feedback design in Sun et al. [41]. Plot (a) shows the average of the maximum error between the converged solution and the theoretical values of each system for both designs. Plot (b) shows the average of the maximum convergence time for each system for both designs.
[…]
The third test was conducted using the feedback circuit proposed by [41] to benchmark the performance of the proposed design. As shown in Figure 16, the proposed equivalent-resistance design is on average twice as fast as the feedback design. In addition, the proposed design reduces the error by a factor of two to four across the systems tested.
IV-D2 Toeplitz Matrix
To benchmark the performance of the proposed design on structured matrices, the same Toeplitz matrix analyzed for the feedback architecture in [42] is reevaluated using the proposed equivalent design. The Toeplitz matrix $T$ for different matrix size $n$ is defined as follows [32]:
$T_{ij}=\frac{1}{\lvert i-j\rvert+1}$ (58)
The time complexity of solving the system $Tx=b$ is governed by the smallest eigenvalue of the matrix $T^{-1}R$, which is upper-bounded by $R_{max}/\lambda_{min}(T)$. For this Toeplitz matrix, $\lambda_{min}(T)$ remains asymptotically constant as $n$ increases, whereas the maximum DD-deviation $R_{max}$ increases linearly with $\log(n)$. Consequently, the expected settling time scales as $O(\log(n))$, matching the complexity reported for the feedback design in [42].
A key distinction from the feedback circuit is that the proposed equivalent design is not dependent on uniform matrix scaling. scaling the matrix by any constant multiplies both $R_{max}$ and $\lambda_{min}(T)$ by the same factor, leaving the ratio, and thus the settling time, unchanged. This property does not hold for feedback architectures.
Figure 17: (a) The minimum eigenvalue of the matrix $T^{-1}R$ for different sizes of the Toeplitz matrix $T$. (b) Settling time for solving the Toeplitz matrices scaled by 1 and 0.1 using the proposed Equivalent design and the Feedback design in [42]. The OAs used for both circuits have 10 MHz GBW.
In practical implementations on memristor-arrays, this Toeplitz matrix require down-scaling because the matrix resistance values fall outside the feasible device range. To evaluate this effect, both $T$ and its scaled version $0.1T$ were solved using the feedback design and the proposed equivalent design for multiple matrix sizes and multiple right-hand-side vectors $b$. The results, summarized in Figure 17, show that the proposed design achieves 2x speedup for the unscaled matrix $T$ and 3x for the scaled matrix $0.1T$ compared to the feedback architecture. These results highlight the scaling resilience of the proposed architecture. This resilience provides more flexibility in the choice of the scaling factor to balance accuracy, power consumption, and hardware constraints without compromising the convergence speed
IV-D3 Covariance Matrix
Figure 18: (a) The 100x100 covariance matrix. (b) SPICE results for the 100 unknowns of solving the covariance matrix. Blue circles are associated with the left axis. The orange triangles are associated with the right axis, showing the relative error for each of the 100 variables.
To evaluate the robustness of the proposed design to circuit nonidealities, specifically device variation and parasitic resistance, the same covariance matrix analyzed for the feedback architecture in [41] is scaled down by a factor of 1000 and reevaluated using the proposed equivalent design.
The unscaled 100x100 covariance matrix $C$ is defined as follows:
[…]
The right-hand-side vector $b$ was randomly generated with values in the range $[0,1]$. All simulations were performed on LTspice v.17.1.8. Using the exact resistance values, the proposed design accurately solved the system with errors below 0.001% compared to the analytical solution as shown in Figure 18.
Figure 19: SPICE simulation results of solving the 100x100 covariance matrix with random resistor device variations within (a) 5% and (b) 10 %.
The first nonideality examined is the device mismatch. In this test, each resistance value of the covariance matrix was randomly varied by $\pm 5\%$ and $\pm 10\%$. The results, shown in Figure 19, indicate that the solution error remained below 12% in both cases except for two points whose analytical values are close to zero. Although the original covariance matrix has a condition number of 15.5, the error does not amplify significantly with device variation. This robustness of the design can be attributed to the fact that the circuit can be described as solving the preconditioned matrix $M^{-1}K$, which is close to the identity matrix for matrices that are nearly diagonal dominant. This lower condition number reduces the error amplification from device variation.
[…]
Figure 21: SPICE simulation results for the transient analysis of thermal noise (bandwidth 1 GHz) of resistors when solving (a) the 100x100 covariance matrix (condition number: 15.6) and (b) a random 100x100 matrix with condition number of 100.
The third nonideality investigated is device noise. In this study, the thermal noise of the variable resistors is modeled as white noise with a bandwidth of 1 GHz. The noise is modelled as a current source parallel to each resistor, using the standard thermal noise model [34]:
$\bar{I_{n}}^{2}=\frac{4kT\Delta f}{R}$ (60)
where $k$ is the Boltzmann constant, $T$ is the absolute temperature (set to 300 K), $R$ is the resistance value and $\Delta f$ is the noise bandwidth. The same transient analysis setup used in the previous studies is employed here, with the supply voltage applied as a step input at 1 $\mu s$. In addition to the covariance matrix, a second 100x100 random SPD system with a condition number of 100 is also analyzed to assess noise sensitivity to condition number. The results, shown in Figure 21, indicate that both systems converge reliably to their respective steady-state solutions despite the presence of thermal noise. In both cases, the resulting fluctuations in the output voltages remain bounded and do not exceed 0.13% of the steady-state values, demonstrating the robustness of the proposed circuit to noise.
The observed robustness to thermal noise can be attributed to the combined effects of damping and implicit preconditioning in the circuit. The implicit preconditioning limits noise amplification, particularly for systems with higher condition number. Meanwhile, the damping ensures that injected white noise appears only as small bounded fluctuations around the steady state solution rather than sustained deviations, guaranteeing stable convergence even under relatively large noise bandwidths.
IV-D4 Operational Amplifier Specifications
@@
The accuracy and speed of the proposed design are influenced by the specifications of the operational amplifiers used. To examine the impact of amplifier choice, the same set of random matrices used in the preliminary design and test 1 of the complexity studies of the spectrum-constrained random systems were simulated using three different amplifiers. The first amplifier, AD712, was used in all previous nonideal simulations. The second amplifier, LTC2050, features a low input offset voltage to minimize errors in the converged solutions. The third amplifier, LTC6268, has relatively high slew rate and gain bandwidth product to reduce settling time. The main specifications of these amplifiers are summarized in Table I.
@@
The simulation results, presented in Figure 22, reveal a tradeoff between speed and accuracy. Using the LTC2050 amplifier reduces the error in the converged solution by an order of magnitude compared to the AD712 but increases the settling time by approximately 1.6 times. Conversely, the LTC6268 amplifier significantly speeds up the system, achieving a two-orders-of-magnitude reduction in settling time, but with an increase in error by an order of magnitude. These highlight that a lower input offset voltage results in higher accuracy in the converged solution, while a higher gain bandwidth product leads to faster system speed.
The selection of the OAs depend on the performance requirements of the target application. For applications such as high-fidelity simulations, where numerical accuracy is prioritized over computational speed, OAs with very low input offset voltage are preferred. On the other hand, for real-time simulation applications, where higher error tolerance is acceptable in exchange for faster convergence, OAs with high gain–bandwidth product are more suitable.
@@ V Discussion
  • Diagonally Dominant SPD Systems: For these systems, the solution is achieved with $O(1)$ complexity, irrespective of the matrix size or specific properties, since the circuit consists exclusively of resistors. This enables the system to reach the theoretical maximum speed.
  • Non-Diagonally Dominant SPD Systems: The accuracy and speed of the solution depend on the specifications of the OAs used and the system matrix properties, particularly, the smallest eigenvalue and the maximum deviation from diagonal dominance. More importantly, the solution speed remains independent of the matrix size.
The reliance on OAs is a widely recognized bottleneck in analog computing architectures, including feedback-based solvers reported in the literature. Power consumption and large area requirements fundamentally limit the feasible number of OAs in a single array, regardless of the specific circuit topology.
The proposed transformation represents a step toward mitigating this bottleneck by reducing the required number of active elements. Instead of scaling with the number of unknowns $n$, the transformation enables the number of active elements to scale with the number of non-diagonally dominant columns $c$. Consequently, an $n\times n$ problem can be solved using at most $4c$ OAs, where $c\ll n$ for matrices that are close to diagonal dominance, such as those arising in finite element solvers for dynamic solid mechanics simulations. The SPICE simulations for systems with a large number of unknowns presented in this work are therefore primarily intended to illustrate convergence and scaling behavior. More advanced matrix transformations that build on the framework introduced here offer a promising path toward further reducing the number of required active elements.
While the proposed design achieves size-independent convergence during the analog solution phase, it requires a preprocessing step to compute the sum of the absolute values of the columns, which has a worst-case digital complexity of $O(n^{2})$. Several approaches can be used to perform this preprocessing, with the choice depending on whether physical area or computational speed is prioritized:
  • System Assembly: In applications such as the Finite Element Method for computational solid mechanics, the calculation of column sums can be naturally incorporated into the system assembly process. During element-wise assembly, the required sums can be accumulated along with the matrix entries with negligible additional overhead. Thus, the preprocessing complexity is still governed by the system assembly complexity rather than the extra column sum operation.
  • Parallel Resistors: Diagonal entries of the submatrix $K_{B}$ can be implemented by representing each column using individual variable resistors connected in parallel, each programmed to the conductance of the corresponding matrix element. This approach eliminates the need for an explicit computational step at the expense of an increase in physical area, which requires $n^{2}$ additional programmable resistive elements.
  • Matrix Vector Multiplication (MVM): An additional analog MVM operation can be used to multiply the matrix by a vector of ones. This operation can be executed using the same crossbar hardware and incurs a constant-time $O(1)$ analog cost.
These alternatives highlight a fundamental design trade-off between circuit area and computational speed, while preserving the size-independent convergence behavior of the proposed analog solver.
@@ VI Conclusion
The presented research work introduces a fundamentally new class of analog compute-in-memory (CIM) solvers, resistive network mapping, that directly maps the given SPD system of equations into a resistive network. This mapping is only realized after leveraging the noninverting OA configuration with some modifications to map negative conductance. Another key aspect of the architecture is the proposed transformation of the original $n\times n$ system (up to $(n^{2}-n)/2$ negative conductance values) into an equivalent $2n\times 2n$ system (up to $n$ negative conductance values). Mapping and solving this transformed system reduces the needed component count by more than 70%, consumes less power, and speeds up the computation by orders of magnitude without additional computational overhead.
@@
For the subclass of symmetric diagonally dominant (SDD) matrices, mapping the transformed system results in a purely passive resistive network whose nodal voltages settle in effectively zero time. This yields true $O(1)$ complexity where the solution speed reaches the theoretical speed limit for analog-CIM architectures and outperforms existing analog solvers, whose settling times are limited by RC constants and feedback loop dynamics, by orders of magnitude. Compared to conventional digital algorithms (even highly optimized sparse solvers), the passive resistive network achieves multiple orders of magnitude speedups for large SDD systems, while consuming a fraction of the energy.
More generally, for any SPD system, the solution speed remains independent of the number of unknowns and depends only on other matrix properties, particularly, the smallest eigenvalue and the maximum deviation from diagonal dominance. The resulting circuit dynamics is shown to correspond to an overdamped second-order differential equation that implicitly solves a preconditioned system, significantly improving numerical conditioning relative to the original problem. Numerical studies on random spectrum-constrained systems confirm the derived complexity bounds, demonstrating that convergence depends on matrix properties rather than the number of unknowns. In addition, the number of active components scales with the number of non–diagonally dominant columns and is therefore always bounded by $n$, further decoupling hardware complexity from problem size.
Benchmarking against the feedback architecture in [41] highlights the advantages of the proposed equivalent design. For both random SPD systems with constrained spectra and structured Toeplitz matrices, the proposed solver consistently achieves faster convergence. Unlike feedback-based designs, the settling time of the proposed circuit is invariant to the uniform scaling of the matrix, providing greater flexibility in selecting resistance ranges to balance accuracy, dynamic range, power consumption, and hardware constraints.
The robustness of the architecture is validated through nonideality analyses. SPICE simulations incorporating realistic OA models, device variations, parasitic resistance, and thermal noise demonstrate stable convergence. The implicit preconditioning inherent to the architecture reduces the sensitivity to component tolerances for matrices with high condition numbers, while the overdamped dynamics suppress noise-induced deviations, ensuring reliable operation under nonideal conditions.
@@
The accuracy of the proposed design is limited by the specifications of the variable resistors and OAs used, which will keep improving with the rapid improvements in device technology. A notable advantage of the proposed design is its compatibility with the cross-point architecture, commonly used in resistive memory arrays. This compatibility facilitates in-memory computations, enabling multiple operations such as MVM within the same framework.
@@
In summary, we show that resistive network mapping has high potential in computational linear algebra by eliminating the mathematical process of solving linear systems of equations. The proposed design offers substantial improvements in speed and efficiency, particularly for SPD systems, such as those arising from Finite Element Analysis in Computational Solid Mechanics. Future work could explore extending this design to a broader range of linear systems and developing more robust solutions for handling non-positive definite matrices efficiently.